# Marginally interpretable spatial logistic regression with bridge processes

### v2

@@ 3.1 Partially collapsed Gibbs sampler
$p(Y,\omega\mid X,\beta,u)$ $=\prod_{i=1}^{n}\prod_{j=1}^{N_{i}}\exp\Big{[}\Big{(}Y_{ij}-\frac{1}{2}\Big{)}\{x_{ij}^{\mathrm{\scriptscriptstyle T}}\beta+u(s_{i})\}-\frac{\omega_{ij}}{2}\{x_{ij}^{\mathrm{\scriptscriptstyle T}}\beta+u(s_{i})\}^{2}\Big{]}\frac{p_{\textsc{pg}}(\omega_{ij})}{2}$ (6)
@@
$(u_{1:n}\mid\lambda)$ $\sim\mathrm{N}_{n}(0,\lambda R),\quad(\lambda\mid\phi)\sim p_{\textsc{m}}(\phi),\quad p(\beta,\phi)=\mathrm{N}_{p}(\beta;\mu_{\beta},\Sigma_{\beta})\times p(\phi)$ (7)
@@
where (6) satisfies $\int p(Y,\omega\mid X,\beta,u)\mathrm{d}\omega=p(Y\mid X,\beta,u)$ corresponding to (1). The log-likelihood conditional on $\omega$ is a quadratic function in both $\beta$ and $u$. Since the spatial random effect $u_{1:n}$ is normal conditional on mixing variable $\lambda$, $u_{1:n}$ can be analytically integrated out, leading to partial collapsing in Step 1 and blocking in Step 2. This leads to improved mixing, especially for the intercept term in $\beta$ that is often highly correlated with $u_{1:n}$. For Step 2(i), we employ particle marginal Metropolis–Hastings (Andrieu et al.,, 2010) to jointly sample $(\phi,\lambda)$ by simulating particles of $\lambda$ from the mixing distribution ${p_{\textsc{m}}}(\phi)$ based on the Theorem 1. Step 2(i) can be further extended to incorporate unknown correlation kernel parameters. The remaining steps and posterior prediction procedure are straightforward, and we defer the detailed derivations to Appendix A.2.
@@ 3.2 Scalable computation with low-rank dependence structure
The Algorithm 1 involves several inversions and determinant calculations of the $n\times n$ matrices, creating a computational bottleneck when the number of spatial locations $n$ is large. Specifically, step 1 involves the inversion of a $n\times n$ matrix and step 2 involves the evaluation of the density and sampling of the multivariate normal distribution of dimension $n$. Without special structures in the corresponding covariance or precision matrices, the computation becomes prohibitive as the number of spatial locations $n$ increases.
Due to the normal mixture representation, several existing computational strategies for Gaussian processes can be easily integrated into bridge processes. One way is to introduce a low-rank structure on the correlation kernel, following the strategy of Finley et al., (2009) for Gaussian processes. Given a positive definite correlation kernel $\mathcal{R}$, denote $R_{qq}=[\mathcal{R}(\tilde{s}_{k},\tilde{s}_{k^{\prime}})]_{k,k^{\prime}=1}^{q}$ as a $q\times q$ matrix formed from $q$ knot locations $\tilde{s}_{k}\in\mathscr{S}$, $k=1,\dots,q$. Then, we consider a correlation kernel $\tilde{\mathcal{R}}$ defined as
~$\tilde{\mathcal{R}}(s,s^{\prime})=r(s)^{\mathrm{\scriptscriptstyle T}}R_{qq}^{-1}r(s^{\prime})+1(s=s^{\prime})\{1-r(s)^{\mathrm{\scriptscriptstyle T}}R_{qq}^{-1}r(s^{\prime})\},$ (8)
where $r(s)=[\mathcal{R}(s,\tilde{s}_{k})]_{k=1}^{q}$ is a $q\times 1$ vector and the term $1(s=s^{\prime})\{1-r(s)^{\mathrm{\scriptscriptstyle T}}R_{qq}^{-1}r(s^{\prime})\}$ in (8) ensures that ~$\tilde{\mathcal{R}}(s,s)=1$. Using the Woodbury matrix identity and the matrix determinant lemma, the computation related to the inversion and calculation of the determinant of $n\times n$ matrices can be reduced to those of $q\times q$ matrices, which gives huge computational benefits when $q\ll n$; see Appendix A.2 for details. In light of Proposition 2, the seamless incorporation of scalable Gaussian process methods through a hierarchical formulation is a distinctive feature of bridge processes.
@@ 4.1 Comparison with existing approaches
First, we perform a simulation study to compare the proposed approach with existing methods for either marginal or conditional inference. As a marginal competitor, we consider the spatial generalized estimating equation method based on pairwise log-odds ratios (Cattelan and Varin,, 2018). As a conditional competitor, we consider a Bayesian logistic mixed model with Gaussian process random effects.
@@
We generate data based on the spatial logistic model (1) under two different processes with bridge-distributed marginals, one with the proposed model and another based on the Gaussian copula process with bridge marginals. This ensures both population-averaged effects $\beta^{M}=\phi\beta$ and site-specific effects $\beta$ are well-defined, where we set $\phi=0.7$. We choose spatial locations uniformly at random from unit square domain $\mathscr{S}=[0,1]^{2}$ to decide training and test location with sizes $(n_{\mathrm{train}},n_{\mathrm{test}})=(200,50)$, where $50$ test locations are held out for assessing predictive performance. We use Matérn correlation kernel $\mathcal{R}_{M}(s,s^{\prime})=(1+\|s-s^{\prime}\|_{2}/\rho)\exp(-\|s-s^{\prime}\|_{2}/\rho)$ with known smoothness $1.5$ and unknown range parameter $\rho$, where true $\rho$ are set as $\rho=0.05$ or $\rho=0.1$. The same correlation kernel is applied to the latent process in the Gaussian copula. We set $N_{i}=10$ for all locations and set $p=2$ including the intercept, where the non-intercept covariate $\{x_{ij}\}$ is generated from an independent standard normal distribution. The true fixed-effect coefficient is set as $\beta=(\beta_{0},\beta_{1})^{\mathrm{\scriptscriptstyle T}}=(0,1)^{\mathrm{\scriptscriptstyle T}}$. This data generation process is repeated 200 times.
@@
We describe details of prior specifications. Following Gelman et al., (2008), we choose Cauchy priors for $\beta$, specifically location 0 scale 10 for $\beta_{0}$ and location 0 scale 1.25 for $\beta_{1}$. For the random effect standard deviation $\sigma_{u}$, we choose a weakly informative half-Cauchy prior (Gelman,, 2006), namely $p(\sigma_{u})=2/\{\pi(1+\sigma_{u}^{2})\}$. For the bridge process models where $\sigma_{u}=3^{-1/2}\pi(\phi^{-2}-1)^{1/2}$, this becomes $p(\phi)=(12)^{1/2}/[\{\pi^{2}-(\pi^{2}-3)\phi^{2}\}(1-\phi^{2})^{1/2}]$ for $\phi\in(0,1)$. For the prior of spatial range $\rho$ in Matérn correlation kernel with known smoothness 1.5, we assign uniform prior $\rho\sim\mathrm{Unif}(0.001,0.3)$ for both bridge and Gaussian process random effect models, which corresponds to effective range approximately between 0.00475 and 1.425 in unit square domain.
@@
Next, we describe inference settings. For the spatial generalized estimating equation method of Cattelan and Varin, (2018) with the same notations therein, we used the number of bins $B=13$, radius $h=0.05$, and $d_{\mathrm{max}}=0.3$ to obtain the empirical spatial lorelogram, and optimize the parameters $\alpha_{2}$ (sill) and $\alpha_{3}$ (range) of Matérn kernel with smoothness 1.5 without nugget. For the bridge and Gaussian process random effect models, we run 11,000 iterations, with the first 1,000 samples discarded as burn-in and 1,000 samples are saved with 10 thin-in rates. The running time of Markov chain Monte Carlo based on a pure R implementation is about 38 mins for the bridge process random effect model and about 20 mins for the Gaussian process random effect model under the Intel(R) Xeon(R) Gold 6336Y 2.40GHz CPU environment.
Table 1: Comparison of population-averaged and site-specific estimates across three methods along with coverage probabilities at nominal level 0.95.
Intercept estimate Fixed effect estimate
$\hat{\beta}_{0}^{M}$ $\hat{\beta}_{0}$ $\hat{\beta}_{1}^{M}$ $\hat{\beta}_{1}$
Data generation SpGEE BrP BrP GP SpGEE BrP BrP GP
Bridge process $\rho=0.05$ Bias 0.001 -0.003 -0.021 -0.022 0.015 0.005 0.004 0.002
RMSE 0.237 0.245 0.413 0.412 0.176 0.128 0.068 0.068
Coverage 91.5% 96.0% 96.0% 96.0% 40.0% 96.5% 96.5% 95.5%
Bridge process $\rho=0.1$ Bias -0.013 -0.002 -0.019 -0.019 0.033 -0.009 0.001 -0.001
RMSE 0.386 0.382 0.621 0.617 0.184 0.140 0.074 0.074
Coverage 85.5% 97.5% 97.5% 97.0% 48.7% 95.0% 94.5% 93.5%
Gaussian copula $\rho=0.05$ Bias -0.005 -0.005 -0.007 -0.005 0.008 0.006 0.001 -0.001
RMSE 0.227 0.237 0.339 0.338 0.063 0.051 0.063 0.063
Coverage 92.0% 96.5% 96.5% 95.5% 94.5% 100.0% 98.5% 98.5%
Gaussian copula $\rho=0.1$ Bias -0.011 0.003 0.008 0.006 0.026 -0.021 0.001 -0.002
RMSE 0.380 0.377 0.568 0.568 0.084 0.069 0.065 0.065
Coverage 85.0% 99.0% 99.0% 98.5% 92.0% 100.0% 98.5% 97.0%

### v3

@@ 3.1 Empirical Bayes approach for attenuation factor
$\mathcal{L}_{\mathrm{comp}}(\bm{\beta},\phi,\rho,\mathbf{u}_{1:n},\lambda)=\mathrm{N}_{n}(\mathbf{u}_{1:n};\bm{0},\lambda\mathbf{R}_{\rho}){p_{\textsc{m}}}(\lambda;\phi)\prod_{i=1}^{n}\left\{\prod_{j=1}^{N_{i}}\frac{\exp(\mathbf{x}_{ij}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}+u(s_{i}))^{y_{ij}}}{1+\exp(\mathbf{x}_{ij}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}+u(s_{i}))}\right\}$ (8)
where $\mathrm{N}_{n}(\mathbf{u}_{1:n};\bm{0},\lambda\mathbf{R}_{\rho})$ is a $n$-dimensional mean zero covariance $\lambda\mathbf{R}_{\rho}$ multivariate normal density evaluated at $\mathbf{u}_{1:n}$. The expression (8) can be viewed as a usual spatial logistic regression with a Gaussian process random effect with further hierarchical formulation on $\lambda$, where mixing density ${p_{\textsc{m}}}(\lambda;\phi)$ is now part of the complete data likelihood. The likelihood $\mathcal{L}(\bm{\beta},\phi,\rho)=\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\mathcal{L}_{\mathrm{comp}}(\bm{\beta},\phi,\rho,\mathbf{u}_{1:n},\lambda)\mathrm{d}\mathbf{u}_{1:n}\mathrm{d}\lambda$ involves high-dimensional integrations which introduces significant computational challenges. We therefore adopt a Bayesian framework, employing Markov chain Monte Carlo for posterior inference on the key parameters of interest, site-specific $\bm{\beta}$ and population-averaged effects ${\bm{\beta}^{\textsc{m}}}=\phi\bm{\beta}$, as well as prediction of probability of binary response at new spatial locations.
However, when observed data is purely spatial and does not have replicates (i.e., with only a single realization $\mathbf{u}_{1:n}$ from the bridge process), fully Bayesian inference for $(\bm{\beta},\phi,\rho)$ faces several challenges. It is well known that in spatial generalized linear models with Gaussian process random effects, the joint estimation of marginal variance and spatial dependence parameters is difficult (Zhang, , 2004). This issue becomes even more pronounced in our setting, as $\phi$ acts as a hyperparameter in $\lambda\sim{p_{\textsc{m}}}(\phi)$ with a deeper hierarchical formulation. While $\phi$ solely controls the marginal variance of the latent bridge process, in the absence of replicates, data provide limited information about $\phi$, resulting in weak identifiability of $\phi$ where the posterior of $\phi$ is highly influenced by its prior (Garrett and Zeger, , 2000).
Although a fully Bayesian approach is possible, we suggest an empirical Bayes procedure that first finds a point estimate $\hat{\phi}$ and then computes the posterior of $(\bm{\beta},\rho)$ given $\phi=\hat{\phi}$. Instead of using marginal maximum likelihood for determining $\hat{\phi}$, which involves high-dimensional integration and is sensitive to model misspecification, we estimate $\hat{\phi}$ by maximizing a composite likelihood based on univariate and bivariate marginals, providing greater robustness (Varin et al., , 2011). Specifically, assuming $N_{i}\geq 2$ for all $i$, we adopt the two-stage composite likelihood method of Zhao and Joe, (2005). In the first stage, leveraging the connection with the marginal logistic regression, we first get a maximum likelihood estimate (MLE) $\hat{\bm{\beta}}^{\textsc{m}}$ from the marginal logistic model (3) ignoring the cluster structure. In the second stage, given $\hat{\bm{\beta}}^{\textsc{m}}$, we set $\hat{\phi}=\operatorname*{arg\,max}_{\phi\in(0,1)}\mathcal{L}_{\mathrm{pair}}(\phi;{\bm{\beta}^{\textsc{m}}}=\hat{\bm{\beta}}^{\textsc{m}})$ using a pairwise composite likelihood constructed from within-cluster pairs, which is
$\mathcal{L}_{\mathrm{pair}}(\phi;{\bm{\beta}^{\textsc{m}}})=\prod_{i=1}^{n}\prod_{j<j^{\prime}}\int_{0}^{\infty}\frac{\exp(\phi^{-1}\mathbf{x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\bm{\beta}^{\textsc{m}}}+u(s_{i}))^{y_{ij}}}{1+\exp(\phi^{-1}\mathbf{x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\bm{\beta}^{\textsc{m}}}+u(s_{i}))}\frac{\exp(\phi^{-1}\mathbf{x}_{ij^{\prime}}^{\mathrm{\scriptscriptstyle T}}{\bm{\beta}^{\textsc{m}}}+u(s_{i}))^{y_{ij^{\prime}}}}{1+\exp(\phi^{-1}\mathbf{x}_{ij^{\prime}}^{\mathrm{\scriptscriptstyle T}}{\bm{\beta}^{\textsc{m}}}+u(s_{i}))}{p_{\textsc{br}}}(u(s_{i});\phi)\mathrm{d}u(s_{i}).$ (9)
It follows from Zhao and Joe, (2005) that, assuming conditional (2) and marginal models (3) are correctly specified but without assumptions on the dependence structure of spatial random effect $u(\cdot)$, the resulting empirical Bayes estimator $\hat{\phi}$ is consistent under increasing-domain asymptotics. This holds because $\hat{\bm{\beta}}^{\textsc{m}}$ is a solution of the unbiased estimating equation arising from the marginal logistic model, and the pairwise likelihood utilizes the bridge-distributed marginal of the random effect but not the spatial dependence information.
3.2 Posterior computation
We describe a two-stage posterior computation procedure in Algorithm 1 based on the empirical Bayes estimate of $\phi$; a description of a fully Bayesian approach is available in Supplementary Section A.2. For simplicity, we focus on prior $p(\bm{\beta},\rho)=p(\bm{\beta})p(\rho)$ with a normal prior for $\bm{\beta}$ and some proper prior on $\rho$.
Stage 1. Find empirical Bayes estimate $\hat{\phi}$ by following steps:
[1] Compute maximum likelihood estimate $\hat{\bm{\beta}}^{\textsc{m}}$ from simple logistic regression (3).
[2] Set $\hat{\phi}=\operatorname*{arg\,max}_{\phi\in(0,1)}\mathcal{L}_{\mathrm{pair}}(\phi;{\bm{\beta}^{\textsc{m}}}=\hat{\bm{\beta}}^{\textsc{m}})$ in eq. (9)
Stage 2. Compute the posterior by MCMC by repeating the following steps:
Repeat:
[1] Sample $\bm{\beta}\sim[\bm{\beta}\mid\lambda^{\mathrm{(old)}},\rho^{\mathrm{(old)}},\bm{\omega}^{\mathrm{(old)}}]$ from multivariate normal.
[2] Sample $\rho\sim[\rho\mid\lambda^{\mathrm{(old)}},\bm{\beta},\bm{\omega}^{\mathrm{(old)}}]$ using adaptive Metropolis-Hastings.
[3] Sample $\lambda\sim[\lambda\mid\rho,\bm{\beta},\bm{\omega}^{\mathrm{(old)}}]$ using independent Metropolis-Hastings
[4] Sample $\mathbf{u}_{1:n}\sim[\mathbf{u}_{1:n}\mid\lambda,\rho,\bm{\beta},\bm{\omega}^{\mathrm{(old)}}]$ from multivariate normal.
[5] Sample $\omega_{ij}\sim[\omega_{ij}\mid\bm{\beta},\mathbf{u}_{1:n}]$ from Pólya-Gamma, independently for all $i,j$.
Algorithm 1 Posterior computation procedure with empirical Bayes estimator for $\phi$.
Stage 2 of Algorithm 1 describes a partially collapsed Gibbs sampler (Van Dyk and Park, , 2008), and we outline key strategies involved. Writing $\mathbf{Y}=\{y_{ij}\}$, $\mathbf{X}=\{\mathbf{x}_{ij}\}$, augmented data $\bm{\omega}=\{\omega_{ij}\}$ and $p_{\textsc{pg}}$ as the density of the Pólya-Gamma(1,0)(1,0) distribution, the complete data model via Pólya-Gamma augmentation (Polson et al., , 2013) and prior becomes
$p(\mathbf{Y},\bm{\omega}\mid\mathbf{X},\bm{\beta},\mathbf{u}_{1:n})$ $=\prod_{i=1}^{n}\prod_{j=1}^{N_{i}}\exp\Big[\Big(y_{ij}-\frac{1}{2}\Big)\{\mathbf{x}_{ij}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}+u(s_{i})\}-\frac{\omega_{ij}}{2}\{\mathbf{x}_{ij}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}+u(s_{i})\}^{2}\Big]\frac{p_{\textsc{pg}}(\omega_{ij})}{2}$ (10)
@@ 3.2 Posterior computation
$(\mathbf{u}_{1:n}\mid\lambda)$ $\sim\mathrm{N}_{n}(0,\lambda\mathbf{R}),\quad\lambda\sim p_{\textsc{m}(\hat{\phi})},\quad p(\bm{\beta},\rho)=\mathrm{N}_{p}(\bm{\beta};\bm{\mu}_{\beta},\bm{\Sigma}_{\beta})\times p(\rho)$ (11)
@@
where (10) satisfies $\int p(\mathbf{Y},\bm{\omega}\mid\mathbf{X},\bm{\beta},\mathbf{u}_{1:n})\mathrm{d}\bm{\omega}=p(\mathbf{Y}\mid\mathbf{X},\bm{\beta},\mathbf{u}_{1:n})$ corresponding to (1). The log-likelihood conditional on $\bm{\omega}$ is a quadratic function in both $\bm{\beta}$ and $\mathbf{u}_{1:n}$. Since the spatial random effect $\mathbf{u}_{1:n}$ is normal conditional on mixing variable $\lambda$, $\mathbf{u}_{1:n}$ can be analytically integrated out, leading to partial collapsing in steps 1 to 3 in Stage 2. This leads to improved mixing, especially for the intercept term in $\bm{\beta}$ that is often highly correlated with $\mathbf{u}_{1:n}$. Conditionally on Pólya-Gamma variables $\bm{\omega}$, we update $\bm{\beta}$ and $\mathbf{u}_{1:n}$ from multivariate normal leveraging conditional conjugacy. When updating $\lambda$, we use Metropolis-Hastings with an independent proposal from the prior. The remaining steps and posterior prediction procedure are straightforward, and we defer the detailed derivations to Supplementary Section A.2.
@@ 3.3 Scalable computation with low-rank dependence structure
The Algorithm 1 involves several inversions and determinant calculations of the $n\times n$ matrices, creating a computational bottleneck when the number of spatial locations $n$ is large. Specifically, step 1 of Stage 2 involves the inversion of an $n\times n$ matrix, and steps 2, 3, and 4 of Stage 2 involve the evaluation of the density and sampling from the multivariate normal distribution of dimension $n$. Without special structures in the corresponding covariance or precision matrices, the computation becomes prohibitive as the number of spatial locations $n$ increases.
Due to the normal mixture representation, several existing computational strategies for Gaussian processes can be easily integrated into bridge processes. One way is to introduce a low-rank structure on the correlation kernel, following the strategy of Finley et al., (2009) for Gaussian processes. Given a positive definite correlation kernel $\mathcal{R}$, denote $\mathbf{R}_{qq}=[\mathcal{R}(\tilde{s}_{k},\tilde{s}_{k^{\prime}})]_{k,k^{\prime}=1}^{q}$ as a $q\times q$ matrix formed from $q$ knot locations $\tilde{s}_{k}\in\mathscr{S}$, $k=1,\dots,q$. Then, we consider a correlation kernel $\tilde{\mathcal{R}}$ defined as
$\tilde{\mathcal{R}}(s,s^{\prime})=\mathbf{r}(s)^{\mathrm{\scriptscriptstyle T}}\mathbf{R}_{qq}^{-1}\mathbf{r}(s^{\prime})+1(s=s^{\prime})\{1-\mathbf{r}(s)^{\mathrm{\scriptscriptstyle T}}\mathbf{R}_{qq}^{-1}\mathbf{r}(s^{\prime})\},$ (12)
where $r(s)=[\mathcal{R}(s,\tilde{s}_{k})]_{k=1}^{q}$ is a $q\times 1$ vector and the term $1(s=s^{\prime})\{1-\mathbf{r}(s)^{\mathrm{\scriptscriptstyle T}}\mathbf{R}_{qq}^{-1}\mathbf{r}(s^{\prime})\}$ in (12) ensures that $\tilde{\mathcal{R}}(s,s)=1$. Using the Woodbury matrix identity and the matrix determinant lemma, the computation related to the inversion and calculation of the determinant of $n\times n$ matrices can be reduced to those of $q\times q$ matrices, which gives huge computational benefits when $q\ll n$; see Supplementary Section A.2 for details. In light of Proposition 3, the seamless incorporation of scalable Gaussian process methods through a hierarchical formulation is a distinctive feature of bridge processes.
One key requirement is that the new correlation kernel $\tilde{R}$ must maintain a unit diagonal $\tilde{\mathcal{R}}(s,s)\equiv 1$. Other possible low-rank constructions include full-scale approximation (Sang and Huang, , 2012); however, ordinary predictive processes (Banerjee et al., , 2008) and nearest-neighbor Gaussian processes (Datta et al., , 2016) are not directly applicable because they induce heterogeneous marginal variances.
@@ 4.1 Comparison with existing approaches
First, we conduct a simulation study to compare the proposed approach with existing methods for both marginal and conditional inference. For the proposed bridge process random effect, we consider both the empirical Bayes approach (BrP) and a fully Bayesian approach that introduces a prior on $\phi$ (BrP-FB). As a marginal effect comparison, we consider the spatial generalized estimating equation (SpGEE) method based on pairwise log-odds ratios (Cattelan and Varin, , 2018). As a conditional effect comparison, we consider a fully Bayesian spatial logistic model with Gaussian process (GP) random effects.
@@
We generate data based on the spatial logistic model (1) under two different processes with bridge-distributed marginals, one with the proposed model and another based on the Gaussian copula process with bridge marginals. This ensures both population-averaged effects ${\bm{\beta}^{\textsc{m}}}=\phi\bm{\beta}$ and site-specific effects $\bm{\beta}$ are well-defined. We consider attenuation factors $\phi\in\{0.7,0.9\}$, where the choice $\phi=0.9$ is motivated by the estimate $\hat{\phi}=0.895$ obtained in the Gambia case study (Section 5). We choose spatial locations uniformly at random from unit square domain $\mathscr{S}=[0,1]^{2}$ to decide training and test locations with sizes $(n_{\mathrm{train}},n_{\mathrm{test}})=(200,50)$, where $50$ test locations are held out for assessing predictive performance. We use Matérn correlation kernel $\mathcal{R}_{M}(s,s^{\prime})=(1+\|s-s^{\prime}\|_{2}/\rho)\exp(-\|s-s^{\prime}\|_{2}/\rho)$ with known smoothness $1.5$ and unknown range parameter $\rho$, where true $\rho$ are set as $\rho=0.05$ or $\rho=0.1$. The same correlation kernel is applied to the latent Gaussian process $\zeta(\cdot)$ in the Gaussian copula. We set $N_{i}=10$ for all locations and set $p=2$ including the intercept, where the non-intercept covariate $\{x_{ij}\}$ is generated from an independent standard normal distribution. The true fixed-effect coefficient is set as $\bm{\beta}=(\beta_{0},\beta_{1})^{\mathrm{\scriptscriptstyle T}}=(0,1)^{\mathrm{\scriptscriptstyle T}}$. This data generation process is repeated 200 times.
@@
We describe details of prior specifications. Following Gelman et al., (2008), we choose Cauchy priors for $\bm{\beta}$, specifically location 0 scale 10 for $\beta_{0}$ and location 0 scale 1.25 for $\beta_{1}$. For the prior of spatial range $\rho$ in Matérn correlation kernel with known smoothness 1.5, we assign a uniform prior $\rho\sim\mathrm{Unif}(0.001,0.3)$ for both bridge and Gaussian process random effect models, which corresponds to an effective range approximately between 0.00475 and 1.425 in unit square domain. For fully Bayesian approaches with prior on the random effect standard deviation $\sigma_{u}$, we choose a weakly informative half-Cauchy prior (Gelman, , 2006), namely $p(\sigma_{u})=2/\{\pi(1+\sigma_{u}^{2})\}$ for Gaussian process random effects. For the bridge process models where $\sigma_{u}=3^{-1/2}\pi(\phi^{-2}-1)^{1/2}$, this becomes $p(\phi)=(12)^{1/2}/[\{\pi^{2}-(\pi^{2}-3)\phi^{2}\}(1-\phi^{2})^{1/2}]$ for $\phi\in(0,1)$, corresponding to prior mean approximately 0.753.
@@
Next, we describe inference settings. For the spatial generalized estimating equation method of Cattelan and Varin, (2018) with the same notations therein, we used the number of bins $B=13$, radius $h=0.05$, and $d_{\mathrm{max}}=0.3$ to obtain the empirical spatial lorelogram, and optimize the parameters $\alpha_{2}$ (sill) and $\alpha_{3}$ (range) of Matérn kernel with smoothness 1.5 without nugget. For the bridge and Gaussian process random effect models, we run 11,000 iterations, with the first 1,000 samples discarded as burn-in and 1,000 samples are saved with 10 thin-in rates. The algorithms are all implemented in R, and the running time of Markov chain Monte Carlo is about 38 mins for the bridge process random effect model and about 20 mins for the Gaussian process random effect model under the Intel(R) Xeon(R) Gold 6336Y 2.40GHz CPU environment.
Table 1: Comparison of estimated population-averaged effects based on 200 simulations in terms of bias, root mean squared error (RMSE), average length of 95% confidence or credible interval (CI.95), and coverage probabilities at nominal level 0.95 (Cover).
Population-averaged ($\hat{\beta}_{1}^{\textsc{m}}$) Data from bridge process Data from Gaussian copula
Setting Method Bias RMSE CI.95 Cover Bias RMSE CI.95 Cover
$\phi=0.7$, $\rho=0.05$ SpGEE 0.015 0.176 0.233 40.0% 0.008 0.063 0.241 94.5%
BrP 0.008 0.174 0.189 33.5% 0.005 0.064 0.190 86.5%
BrP-FB 0.005 0.128 0.532 96.5% 0.006 0.051 0.568 100.0%
$\phi=0.7$, $\rho=0.1$ SpGEE 0.033 0.184 0.275 48.5% 0.026 0.084 0.293 92.0%
BrP 0.030 0.182 0.193 31.0% 0.021 0.084 0.192 75.5%
BrP-FB -0.009 0.140 0.557 95.0% -0.021 0.069 0.608 100.0%
$\phi=0.9$, $\rho=0.05$ SpGEE -0.033 0.151 0.230 70.5% 0.002 0.058 0.228 95.0%
BrP -0.038 0.151 0.212 69.5% -0.003 0.057 0.218 94.0%
BrP-FB -0.087 0.139 0.429 95.5% -0.099 0.114 0.416 95.0%
$\phi=0.9$, $\rho=0.1$ SpGEE -0.030 0.148 0.252 74.0% 0.010 0.069 0.244 90.5%
BrP -0.036 0.151 0.212 69.0% 0.005 0.068 0.219 88.0%
BrP-FB -0.105 0.156 0.464 94.0% -0.117 0.137 0.470 94.5%